OpenStudy (eyka94):

Find the limit please! lim[x:1,(x^4-1)/ (x^3-1)]

OpenStudy (anonymous):

factor and cancel

OpenStudy (eyka94):

@satellite73 How to factor ? :(

OpenStudy (anonymous):

\[x^4-1=(x+1)(x-1)(x^2+1)\] \[x^3-1=(x-1)(x^2+x+1)\]

OpenStudy (eyka94):

Thanks. Isnt that a basic algebra I didn't learn that in school :|

OpenStudy (anonymous):

top one is difference of two squares bottom is difference of two cubes

OpenStudy (anonymous):

@eyka94 even if you do not know the algebra, you know that if you replace \(x\) by \(1\) you get zero top and bottom that means you can factor both as \((x-1)(\text{something})\) you can find the something by division if you have to

OpenStudy (eyka94):

@satellite73 Thank you so much ;)

OpenStudy (anonymous):

yw